Conrad, Inc. recently lost a portion of its records in an office fire. The following information was salvaged from the accounting records.

Cost of Goods Sold $ 69,500
Work-in-Process Inventory, Beginning 12,300
Work-in-Process Inventory, Ending 9,900
Selling and Administrative Expense 17,250
Finished Goods Inventory, Ending 17,475
Finished Goods Inventory, Beginning ?
Direct Materials Used ?
Factory Overhead Applied 12,900
Operating Income 14,495
Direct Materials Inventory, Beginning 11,405
Direct Materials Inventory, Ending 6,315
Cost of Goods Manufactured 64,230

Direct labor cost incurred during the period amounted to 1.5 times the factory overhead. The CFO of Conrad, Inc. has asked you to recalculate the following accounts and to report to him by the end of the day.

What is the amount of net sales?
Answer #1

Answer:-

Net sales = 101,245

Purchased direct materials = 24490

Finished goods inventory (beginning)= 22745

As per the given information explanation is given below :

To calculate the cost of manufactured goods we need to use the following formula:

Cost of good manufactured= Beginning work in progress+ (beginning inventory direct material + purchase direct material - ending inventory direct material) + direct labor + manufactured overhead - ending work in progress

64230=12300+(11405 +purchase direct material -6315)+19350+12900-9900

Purchase direct material = 64230-12300-11405+6315-19350-12900+9900

Purchase direct material = 24490

As we know :

COGS= Beginning Inventory+Production during period−Ending Inventory

69500= Beginning Inventory + 64230-17475

Beginning Inventory = 69500+17475 -64230

Beginning Inventory= 22745

NET SALES = COST OF GOODS SOLD + OPERATING INCOME + SELLING AND ADMINISTRATIVE EXPENSES

NET SALES = 69500+14495+17250

NET SALES =101,245

* DIRECT LABOUE IS GIVEN IN THE QUESTION AS 1.5 OF Factory Overhead

HENCE IT WILL BE =12900*1.5 = 19350